xylose is a carbohydrate that contains carbon, hydrogen, and oxygen. The empirical formula of Xylose is CH2O. and its molar mass 150.1 g/mol. Find the molecular formula of xylose
Answer #1

Molar mass of CH2O,

MM = 1*MM(C) + 2*MM(H) + 1*MM(O)

= 1*12.01 + 2*1.008 + 1*16.0

= 30.026 g/mol

Now we have:

Molar mass = 150.1 g/mol

Empirical formula mass = 30.026 g/mol

Multiplying factor = molar mass / empirical formula mass

= 150.1/30.026

= 5

So molecular formula is:C5H10O5

Answer: C5H10O5